# Amortized SHAP values via sparse Fourier function approximation

### v1

@@ 3.2 Sparsity of ensembles of decision trees
Thus, the Fourier transform of a decision tree can be computed recursively (Kushilevitz & Mansour, 1993; Mansour, 1994). The degree of a decision tree function of depth $d$ is $d$, and if $|\text{supp}(t_{\mathrm{left}})|=k_{\mathrm{left}}$ and $|\text{supp}(t_{\mathrm{right}})|=k_{\mathrm{right}}$, then $|\text{supp}(t)|\leq 2(k_{\mathrm{left}}+k_{\mathrm{right}})$. As a result, a decision tree function is $k$-sparse, where $k=O(4^{d})$, although in some cases, when the decision tree is not balanced or cancellations occur, the Fourier transform can be sparser, i.e., admit a lower $k$, than the above upper bound on the sparsity suggests.
Due to the linearity of the Fourier transform, the Fourier transform of an ensemble of trees, such as those produced by the random forest, cat-boost (Dorogush et al., 2018), and XGBoost (Chen & Guestrin, 2016) algorithms/libraries, can be computed by taking the average of the Fourier transform of each tree. If the random forest model has $T$ trees, then its Fourier transform is $k=O(T4^{d})$-sparse and of degree $d$ equal to its maximum depth of the constituent trees.
@@ 4 Computing SHAP values with Fourier representation of functions
In the previous section, we saw that many real-world models trained on tabular/discrete data can be exactly represented (in the case of ensembles of decision trees), or efficiently approximated (in the case of neural networks) using a compact (sparse) Fourier representation. We saw neural networks have a tendency to learn approximately low-degree, and hence by Proposition 2 sparse, functions. This has been attributed in numerous works to the reason why they generalize well and do not overfit despite their over-parameterized nature (Valle-Perez et al., 2018; Yang & Salman, 2019; Huh et al., 2022; Durvasula & Liter, 2020; Kalimeris et al., 2019; Neyshabur et al., 2017; Arpit et al., 2017). We also saw that (ensembles) of decision trees, by nature, have sparse Fourier representations (Kushilevitz & Mansour, 1993; Mansour, 1994). More generally, as made formal in Proposition 1 and the remarks after, any “simple” function that can be written as a summation of a “few” functions each depending on a “few” of the input variables is sparse and low-degree in the Fourier domain.
We propose the following method to approximate SHAP values for black-box functions. In the first step, given query access to a black-box function, we utilize a sparse Fourier approximation algorithm such as (Li & Ramchandran, 2015; Amrollahi et al., 2019; Li et al., 2015) to efficiently extract its sparse and hence compactly represented Fourier approximation. See Appendix B.2 for a more detailed explanation. Next, in our second step presented here, we use the Fourier representation to exactly compute SHAP values.
Let $h:\{0,1\}^{n}\rightarrow\mathbb{R}$ be some predictor which we assume to be $k$-sparse with $n$ binary input features. Let $[n]\triangleq\{1,\dots,n\}$ be the set of features and let $x^{*}\in\{0,1\}^{n}$ be the instance we are explaining. As outlined in Equation 1, SHAP values are obtained from the following equation:
$\phi_{i}^{h}=\sum\limits_{S\subseteq[n]\setminus\{i\}}\frac{|S|!(n-|S|-1)!}{M!}(v_{h}(S\cup\{i\})-v_{h}(S)),$ $\forall i\in[n]=\{1,2,\dots,n\}$
where $v_{h}(S)$ is the average prediction when one only knows $x_{S}^{*}$. The average is taken over the (background) dataset $\mathcal{D}=\{(x,y)^{i}\}$ (Lundberg & Lee, 2017). More precisely:
$v_{h}(S)\triangleq\frac{1}{|\mathcal{D}|}\sum\limits_{(x,y)\in\mathcal{D}}h(x^{*}_{S},x_{[n]\setminus S})$
The above equations show the SHAP values are linear with respect to the prediction function $h$. Therefore we proceed by computing the Shapley values for a single Fourier basis function $\Psi_{f}(x)=(-1)^{\langle f,x\rangle},f\in\{0,1\}^{n}$:
@@
This expression still has an exponential (in $n$) sum, since we are summing over all subsets $S$. As a main contribution, we find a closed-form analytic expression for the inner summation using a combinatorial argument. This results in the following key Lemma:
@@
where $A\triangleq\{j\in[n]|x_{j}\neq x^{*}_{j},j\neq i,f_{j}=1\}$. (Proof in Appendix C.2)
@@
where $A$ is the same as in Lemma 1. (Proof in Appendix C.3)
Theorem 2 gives us a computationally efficient way to go from a Fourier representation of a function $h$ to SHAP values. The SHAP values from this equation are exact, i.e., as long as the Fourier representation is exact, the SHAP values are also precise values. This is in contrast to KernelSHAP, where the SHAP values are approximated using stochastic sampling and one needs to check for convergence to make sure the approximation is accurate. The approximation in our method is constrained to the first step: computing the (approximate) sparse Fourier representation of the black-box.
@@
We implement our algorithm called FourierShap using JAX (Bradbury et al., 2018), which allows for fast vectorized operations on GPUs. Each term inside the summations of Equation 5 can be implemented with simple vector operations. Furthermore, summations over the $k$ different frequencies in the support of $h$ and also background data points can both be efficiently implemented and parallelized using this library using its vmap operator. We perform all upcoming experiments on a single GPU. Nevertheless, we believe faster computation can also be achieved by crafting dedicated code designed to efficiently compute Equation 5 on GPU.
@@ 5 Experiments
We assess the performance of our algorithm, FourierSHAP, on four different real-world datasets of varying nature and dimensionality. Three of our datasets are related to protein fitness landscapes (Poelwijk et al., 2019; Wu et al., 2016; Sarkisyan et al., 2016) and are referred to as “Entacmaea” (dimension $n=13$), “GB1” ($n=80$), and “avGFP” ($n=236$) respectively. The fourth dataset is a GPU-tuning (Nugteren & Codreanu, 2015) dataset referred to as “SGEMM” ($n=40$). The features of these datasets are binary (zero-one) and/or categorical with standard one-hot encodings. See Appendix D for dataset details.
For the Entacmaea and SGEMM datasets, we train fully connected neural networks with 3 hidden layers containing 300 neurons each. For GB1 we train ensembles of trees models of varying depths using the random forest algorithm and for avGFP we train again, ensembles of trees models of varying depths using the cat-boost algorithm/library (Dorogush et al., 2018).
Black-box setting. The first step of the FourierSHAP algorithm is to compute a sparse Fourier approximation of the black-box model. We use a GPU implementation of a sparse Walsh-Hadamard Transform (sparse WHT) a.k.a Fourier transform algorithm (Amrollahi et al., 2019) for each of the four trained models. The algorithm accepts a sparsity parameter $k$ which is the sparsity of the computed Fourier representation. Higher sparsity parameter $k$ results in a better function approximation but the sparse-WHT runtime increase linearly in $k$ as well. In Figure 1 we plot the accuracy of the Fourier function approximation as measured by the $R^{2}$-score for different values of $k$ (which result in different runtimes). The $R^{2}$ score is computed over a dataset formed by randomly sampling the Boolean cube $\{0,1\}^{n}$.
@@
The second step of FourierSHAP utilizes the Fourier approximation to compute SHAP values using Equation 5. We implement this step using JAX library (Bradbury et al., 2018), and run it on a GPU. For each model to be explained, we choose four different values for the number of background samples and four different values for the number of query points to be explained, resulting in a total of 16 runs of FourierSHAP for each model. Error bars capture these variations. We take the runtime of KernelSHAP, with Github repo default settings, to be the base runtime all other methods are compared to, i.e., we assume its runtime is 1 unit.
Figure 2: Speedup vs. Accuracy. Speedup of different algorithms is reported as a multiple compared to the runtime of KernelShap. Accuracy is quantified by the $R^{2}$-score against ground truth SHAP values. DeepLift(Shrikumar et al., 2017): is a white-box algorithm for neural networks. LinRegShap(Covert & Lee, 2020) is black-box algorithm and a variance-reduced version of KernelShap. FastShap(Jethani et al., 2021), a black box algorithm, is a trained MLP to output SHAP values given inputs in one forward pass. We experiment three sizes for MLPs for each dataset. FourierShap is ours. We are 10-10000x faster than LinRegShap on all dataset/model variations. More notably, we outperform DeepLift (a white-box algorithm) in the neural network model setting even though we assume only query access (black-box setting) to the neural network. We achieve higher accuracy than FastShap in 3/4 settings, while enabling a fine-grained control over the speed-accuracy trade-off.
In order to measure the accuracy of the SHAP values produced by our and other algorithms we need ground truth SHAP values. We use KernelSHAP to this extent. Note that KernelSHAP is inherently an approximation method for computing interventional SHAP values. However, this approximation becomes more precise by sampling more coalition subsets. Therefore, to generate ground truth values, for each dataset, we sample more and more and check for convergence in these values. A more detailed explanation can be found in Appendix F.3.
@@
We compare against the following baselines in Figure 2. The first is LinRegShap, a variance-reduced version of KernelSHAP (Covert & Lee, 2020). We found that although this algorithm requires fewer queries from the black-box, it takes orders of magnitudes longer to run compared to ours. Secondly, for the neural network models, we also compare against a state-of-the-art white-box method – DeepLift (Shrikumar et al., 2017). This algorithm, requires access to the neural network’s activations in all layers. In comparison, we achieve a 10-100x speedup while being both more accurate and only assuming query access to the neural net (true black-box setting).
@@
Finally, we compare against FastShap (Jethani et al., 2021), which is the closest to our setting, i.e., the model agnostic black-box, amortized setting. We provide higher accuracy for similar values of speedup compared to FastShap in 3/4 settings. More importantly though, we can see by controlling the sparsity parameter $k$ of the Fourier function approximation, we can control the tradeoff between accuracy and speed in a reliable and fine-grained manner. Whereas for FastShap the accuracy of the SHAP values relies on the functional approximation properties of the MLP which directly produces SHAP values. As seen in Figure 2, the MLP can behave in unpredictable ways. In this figure we can see that increasing the depth of the model (and hence the approximation capability of the MLP) does not have a reliable effect on SHAP accuracy. Finally, in FastShap for a different choice of background dataset a new MLP model has to be trained from scratch. Since, the MLP approximates the process of directly computing SHAP values. Whereas we can support different sets of background datasets since our Fourier functional approximation is done on the black box predictor, and not the whole process of producing SHAP values in an end-to-end manner (FastShap).
@@
Tree setting. FourierSHAP can also be utilized for the computation of SHAP values for (ensembles of) trees in the white-box setting, i.e., where full access to the tree’s structure is available. In this setting, the exact sparse Fourier representation of an ensemble of trees can be efficiently computed (the first step of FourierSHAP) using Equation 2. With the exact Fourier representation at hand, SHAP values can be efficiently and exactly computed using Equation 5, the second step of FourierSHAP. This makes our method a highly parallelizable alternative for TreeShap (Lundberg et al., 2020) with the potential for order of magnitude speedups.
@@
To demonstrate our method’s ability in fast computation of SHAP values in this setting, we compute SHAP values for random forests fitted on all aforementioned real-world datasets. We compare FourierSHAP’s speedup over TreeSHAP (Lundberg et al., 2020), to FastTreeSHAP (Yang, 2021), a fast implementation of TreeSHAP, and the GPU implementation of TreeSHAP (Mitchell et al., 2022b). To the best of our knowledge, these are the fastest available frameworks for computation of the “interventional” SHAP values. Figure 3 shows the superior speed of our method over all state-of-the-art algorithms in most settings, which weakens by the increment of depth and number of features, resulting in more frequencies and computation overhead. Note that the SHAP values computed by all methods are precise and identical to the values produced by TreeSHAP, which is to be expected for FourierSHAP, given access to exact Fourier representation of random forests.
@@
Figure 3: Speedup vs. depth of tree, for different algorithms, reported as a multiple compared to the runtime of TreeSHAP Lundberg et al. (2020). FourierSHAP is ours. As other baselines we have a GPU implementation of TreeSHAP (Mitchell et al., 2022b), and FastTreeSHAP (Yang & Salman, 2019) . We achieve order of magnitude speedups over all depths on the Entacmaea and SGEMM datasets. We also achieve significant speedups in the other two datasets; however, the edge diminishes as the maximum depth increases.
@@ Conclusions
Appendix A Appendix
Appendix B Relevant work
B.1 Neural network theory and simplicity biases
With regards to simplicity biases in fully connected neural networks, a substantial amount of research has been dedicated to analyzing neural networks in function space. This line of research is dedicated to firstly showing that “infinite-width”, randomly initialized (with Gaussian distribution) neural networks are distributed as Gaussian Processes (GPs) and secondly computing the kernel associated to the GP (Neal, 1996; Williams, 1996; Cho & Saul, 2009; Hazan & Jaakkola, 2015; Lee et al., 2017; Ancona et al., 2018; Daniely, 2017). The kernel associated with the GP is commonly known as the “conjugate kernel” (Daniely, 2017) or the “NN-GP kernel”(Lee et al., 2017). Other works show that in infinite-width neural networks weights after training via SGD do not end up too far from the original (Chizat et al., 2019; Jacot et al., 2018a; Du et al., 2018; Allen-Zhu et al., 2019a; b), referred to as “lazy training” by Chizat et al. (2019). This allowed Jacot et al. (2018a) to prove that the evolution of an infinite-width neural network during training can be described by a kernel called the “Neural Tangent Kernel”. Lee et al. (2019) showed, through extensive experiments that the same behavior holds even for the more realistic case of neural nets of finite width. Empirically speaking, it was shown by Rahaman et al. (2019) that a neural net with one input tends to learn sinusoids of lower frequencies in earlier epochs than those with higher frequencies. By analyzing the spectrum of the NTK’s Gram matrix, Ronen et al. (2019); Basri et al. (2020) were able to formally prove this empirical finding. Yang & Salman (2019); Fan & Wang (2020) analyze the spectra of the NTK gram matrix for higher dimensional inputs. Specifically, Yang & Salman (2019); Valle-Perez et al. (2018) provide simplicity bias results for the case where the inputs to the neural net are Boolean (zero-one) vectors.
B.2 Sparse and low-degree Fourier transform algorithms

### v2

@@ 3.2 Sparsity of ensembles of decision trees
Thus, the Fourier transform of a decision tree can be computed recursively [Kushilevitz and Mansour, 1993, Mansour, 1994]. The degree of a decision tree function of depth $d$ is $d$, and if $|\text{supp}(t_{\mathrm{left}})|=k_{\mathrm{left}}$ and $|\text{supp}(t_{\mathrm{right}})|=k_{\mathrm{right}}$, then $|\text{supp}(t)|\leq 2(k_{\mathrm{left}}+k_{\mathrm{right}})$. As a result, a decision tree function is $k$-sparse, where $k=O(4^{d})$, although in some cases, when the decision tree is not balanced or cancellations occur, the Fourier transform can be sparser, i.e., admit a lower $k$, than the above upper bound on the sparsity suggests.
Due to the linearity of the Fourier transform, the Fourier transform of an ensemble of trees, such as those produced by the random forest, cat-boost [Dorogush et al., 2018], and XGBoost [Chen and Guestrin, 2016] algorithms/libraries, can be computed by taking the average of the Fourier transform of each tree. If the random forest model has $T$ trees, then its Fourier transform is $k=O(T4^{d})$-sparse and of degree $d$ equal to its maximum depth of the constituent trees.
@@ 4 Computing SHAP values with Fourier representation of functions
In the previous section, we saw that many real-world models trained on tabular/discrete data can be exactly represented (in the case of ensembles of decision trees), or efficiently approximated (in the case of neural networks) using a compact (sparse) Fourier representation. We saw neural networks have a tendency to learn approximately low-degree, and hence by Proposition 2, sparse functions. This has been attributed in numerous works to the reason why they generalize well and do not overfit despite their over-parameterized nature [Valle-Perez et al., 2018, Yang and Salman, 2019, Huh et al., 2022, Durvasula and Liter, 2020, Kalimeris et al., 2019, Neyshabur et al., 2017, Arpit et al., 2017]. We also saw that (ensembles) of decision trees, by nature, have sparse Fourier representations [Kushilevitz and Mansour, 1993, Mansour, 1994]. More generally, as made formal in Proposition 1 and the remarks after, any “simple” function that can be written as a summation of a “few” functions each depending on a “few” of the input variables is sparse and low-degree in the Fourier domain.
We propose the following method to approximate SHAP values for black-box functions. In the first step, given query access to a black-box function, we utilize a sparse Fourier approximation algorithm such as [Li and Ramchandran, 2015, Amrollahi et al., 2019, Li et al., 2015] to efficiently extract its sparse and hence compactly represented Fourier approximation. See Appendix A.2 for a more detailed explanation. Next, in our second step presented here, we use the Fourier representation to exactly compute SHAP values.
Let $h:\{0,1\}^{n}\rightarrow\mathbb{R}$ be some predictor that we assume to be $k$-sparse. Assume $[n]\triangleq\{1,\dots,n\}$ as the set of features and let $x^{*}\in\{0,1\}^{n}$ be the instance we are explaining. Aligned with [Lundberg and Lee, 2017], we use $v_{h}(S)\triangleq\frac{1}{|\mathcal{D}|}\sum\limits_{(x,y)\in\mathcal{D}}h(x^{*}_{S},x_{[n]\setminus S})$ as the value function, which is the average prediction when fixing $x_{S}^{*}$ in the (background) dataset $\mathcal{D}$. We compute the Shapley values for a single Fourier basis function $\Psi_{f}(x)=(-1)^{\langle f,x\rangle},f\in\{0,1\}^{n}$ according to Equation 1:
@@
This expression still has an exponential sum (in $n$), since we are summing over all subsets $S$. As a main contribution, we find a closed-form analytic expression for the inner summation using a combinatorial argument. This results in the following key Lemma:
@@
where $A\triangleq\{j\in[n]|x_{j}\neq x^{*}_{j},j\neq i,f_{j}=1\}$. (Proof in Appendix B.2)
@@
where $A$ is the same as in Lemma 1. (Proof in Appendix B.3)
Theorem 2 gives us a computationally efficient way to go from a Fourier representation of a function $h$ to SHAP values. The SHAP values from this equation are exact, i.e., as long as the Fourier representation is exact, the SHAP values are also precise values. This is in contrast to KernelSHAP, where the SHAP values are approximated using stochastic sampling and one needs to check for convergence to ensure the accuracy. The approximation in our method is constrained to computing the sparse Fourier representation, only in the black-box setting.
@@
We implement our algorithm called FourierShap using JAX [Bradbury et al., 2018], which allows for fast vectorized operations on GPUs. Each term inside the summations of Equation 5 can be implemented with simple vector operations. Furthermore, summations over the $k$ different frequencies in the support of $h$ and also background data points can both be efficiently implemented and parallelized using this library using its vmap operator. We perform all upcoming experiments on a single GPU. Nevertheless, we believe faster computation can also be achieved by crafting dedicated code designed to efficiently compute Equation 5 on GPU.
@@ 5 Experiments
We assess the performance of our algorithm, FourierSHAP, on four different real-world datasets of varying nature and dimensionality. Three of our datasets are related to protein fitness landscapes [Poelwijk et al., 2019, Wu et al., 2016, Sarkisyan et al., 2016] and are referred to as “Entacmaea” (dimension $n=13$), “GB1” ($n=80$), and “avGFP” ($n=236$) respectively. The fourth dataset is a GPU-tuning [Nugteren and Codreanu, 2015] dataset referred to as “SGEMM” ($n=40$). The features of these datasets are binary (zero-one) and/or categorical with standard one-hot encodings. See Appendix C for dataset details.
For the Entacmaea and SGEMM datasets, we train fully connected neural networks with 3 hidden layers containing 300 neurons each. For GB1 we train ensembles of trees models of varying depths using the random forest algorithm and for avGFP we train again, ensembles of trees models of varying depths using the cat-boost algorithm/library [Dorogush et al., 2018].
Black-box setting. The first step of the FourierSHAP algorithm is to compute a sparse Fourier approximation of the black-box model. We use a GPU implementation of a sparse Walsh-Hadamard Transform (sparse WHT) a.k.a Fourier transform algorithm [Amrollahi et al., 2019] for each of the four trained models. The algorithm accepts a sparsity parameter $k$ which is the sparsity of the computed Fourier representation. Higher sparsity parameter $k$ results in a better function approximation but the sparse-WHT runtime increase linearly in $k$ as well. In Figure 1 we plot the accuracy of the Fourier function approximation as measured by the $R^{2}$-score for different values of $k$ (which result in different runtimes). The $R^{2}$ score is computed over a dataset formed by randomly sampling the Boolean cube $\{0,1\}^{n}$.
@@
The second step of FourierSHAP utilizes the Fourier approximation to compute SHAP values using Equation 5. We implement this step using JAX library [Bradbury et al., 2018], and run it on a GPU. For each model to be explained, we choose four different values for the number of background samples and four different values for the number of query points to be explained, resulting in a total of 16 runs of FourierSHAP for each model. Error bars capture these variations. We take the runtime of KernelSHAP, with Github repo default settings, to be the base runtime all other methods are compared to, i.e., we assume its runtime is 1 unit.
Figure 2: Speedup vs. Accuracy. Speedup of different algorithms is reported as a multiple compared to the runtime of KernelShap. Accuracy is quantified by the $R^{2}$-score against ground truth SHAP values. DeepLift is a white-box algorithm for neural networks. LinRegShap is a black-box algorithm and a variance-reduced version of KernelShap. FastShap, a black-box algorithm, is a trained MLP to output SHAP values given inputs in one forward pass. We test MLPs of three different sizes for FastShap. FourierShap is ours. We are 10-10000x faster than LinRegShap on all dataset/model variations. More notably, we outperform DeepLift in the neural network model setting even though we assume only query access (black-box setting). We achieve higher accuracy than FastShap in 3/4 settings, while enabling a fine-grained control over the speed-accuracy trade-off.
In order to measure the accuracy of the SHAP values produced by our and other algorithms we need ground truth SHAP values. We use KernelSHAP to this extent. Note that KernelSHAP is inherently an approximation method for computing interventional SHAP values. However, this approximation becomes more precise by sampling more coalition subsets. Therefore, to generate ground truth values, for each dataset, we sample more and more coalition subsets and check for convergence in these values. A more detailed explanation can be found in Appendix E.3.
@@
We compare against the following baselines in Figure 2. The first is LinRegShap, a variance-reduced version of KernelSHAP [Covert and Lee, 2020]. We found that although this algorithm requires fewer queries from the black-box, it takes orders of magnitudes longer to run compared to ours. Secondly, for the neural network models, we also compare against a state-of-the-art white-box method – DeepLift [Shrikumar et al., 2017]. This algorithm, requires access to the neural network’s activations in all layers. In comparison, we achieve a 10-100x speedup while being both more accurate and only assuming query access to the neural net (true black-box setting).
@@
Finally, we compare against FastShap [Jethani et al., 2021], which is the closest to our setting, i.e., the model agnostic black-box, amortized setting. We provide higher accuracy for similar values of speedup compared to FastShap in 3/4 settings. More importantly though, we can see by controlling the sparsity parameter $k$ of the Fourier function approximation, we can control the tradeoff between accuracy and speed in a reliable and fine-grained manner. Whereas for FastShap the accuracy of the SHAP values relies on the functional approximation properties of the MLP which directly produces SHAP values. As seen in Figure 2, the MLP can behave in unpredictable ways. In this figure we can see that increasing the depth of the model (and hence the approximation capability of the MLP) does not have a reliable effect on SHAP accuracy. Finally, in FastShap for a different choice of background dataset a new MLP model has to be trained from scratch. Since, the MLP approximates the process of directly computing SHAP values. Whereas we can support different sets of background datasets since our Fourier functional approximation is done on the black box predictor, and not the whole process of producing SHAP values in an end-to-end manner (FastShap).
@@
Tree setting. FourierSHAP can also be utilized for the computation of SHAP values for (ensembles of) trees in the white-box setting, i.e., where full access to the tree’s structure is available. In this setting, the exact sparse Fourier representation of an ensemble of trees can be efficiently computed (the first step of FourierSHAP) using Equation 2. With the exact Fourier representation at hand, SHAP values can be efficiently and exactly computed using Equation 5, the second step of FourierSHAP. This makes our method a highly parallelizable alternative for TreeShap [Lundberg et al., 2020] with the potential for order of magnitude speedups.
@@
To demonstrate our method’s ability in fast computation of SHAP values in this setting, we compute SHAP values for random forests fitted on all aforementioned real-world datasets. We compare FourierSHAP’s speedup over TreeSHAP [Lundberg et al., 2020], to FastTreeSHAP [Yang, 2021], a fast implementation of TreeSHAP, the GPU implementation of TreeSHAP [Mitchell et al., 2022b] and PLTreeSHAP [Zern et al., 2023]. To the best of our knowledge, these are the fastest available frameworks for computation of the “interventional” SHAP values. Figure 3 shows the superior speed of our method over all state-of-the-art algorithms in most settings, which weakens by the increment of depth and number of features, resulting in more frequencies and computation overhead. Note that the SHAP values computed by all methods are precise and identical to the values produced by TreeSHAP, which is to be expected for FourierSHAP, given access to exact Fourier representation of random forests.
@@
Figure 3: Speedup vs. depth of tree, for different algorithms, reported as a multiple compared to the runtime of TreeSHAP. FourierSHAP is ours. As other baselines we have a GPU implementation of TreeSHAP, FastTreeSHAP, and PLTreeSHAP. We achieve order of magnitude speedups over all depths on the Entacmaea and SGEMM datasets. We also achieve significant speedups in the other two datasets; however, the edge diminishes as the maximum depth increases.
@@ Conclusions
Appendix A Relevant work
A.1 Neural network theory and simplicity biases
With regards to simplicity biases in fully connected neural networks, a substantial amount of research has been dedicated to analyzing neural networks in function space. This line of research is dedicated to firstly showing that “infinite-width”, randomly initialized (with Gaussian distribution) neural networks are distributed as Gaussian Processes (GPs) and secondly computing the kernel associated to the GP [Neal, 1996, Williams, 1996, Cho and Saul, 2009, Hazan and Jaakkola, 2015, Lee et al., 2017, Ancona et al., 2018, Daniely, 2017]. The kernel associated with the GP is commonly known as the “conjugate kernel” [Daniely, 2017] or the “NN-GP kernel”[Lee et al., 2017]. Other works show that in infinite-width neural networks weights after training via SGD do not end up too far from the original [Chizat et al., 2019, Jacot et al., 2018a, Du et al., 2018, Allen-Zhu et al., 2019a, b], referred to as “lazy training” by Chizat et al. [2019]. This allowed Jacot et al. [2018a] to prove that the evolution of an infinite-width neural network during training can be described by a kernel called the “Neural Tangent Kernel”. Lee et al. [2019] showed, through extensive experiments that the same behavior holds even for the more realistic case of neural nets of finite width. Empirically speaking, it was shown by Rahaman et al. [2019] that a neural net with one input tends to learn sinusoids of lower frequencies in earlier epochs than those with higher frequencies. By analyzing the spectrum of the NTK’s Gram matrix, Ronen et al. [2019], Basri et al. [2020] were able to formally prove this empirical finding. Yang and Salman [2019], Fan and Wang [2020] analyze the spectra of the NTK gram matrix for higher dimensional inputs. Specifically, Yang and Salman [2019], Valle-Perez et al. [2018] provide simplicity bias results for the case where the inputs to the neural net are Boolean (zero-one) vectors.
A.2 Sparse and low-degree Fourier transform algorithms